OpenStudy (anonymous):

What is the slope of the line 5x - 2y = -10?

OpenStudy (cwrw238):

rearrange the equation to the form y = mx + c - your slope will be value of m

OpenStudy (anonymous):

i know, but i don't know how to do that

OpenStudy (anonymous):

I think there is no need to go this way here..

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

When you are given with the equation in the form : \(ax + by = c\) Then slope is given by: \[Slope = -\frac{x \; \; coefficient}{y \; \; coefficient}\]

OpenStudy (anonymous):

Hey, you would first subtract 5x from both sides. SO you would get -2y=-5x -10 Then you would divide both sides to isolate the "y". so, y=5/2x +5:) and slope=m in the equation y=mx+b:) Hope I helped!

OpenStudy (anonymous):

Thank you soooo much!

OpenStudy (anonymous):

Here x coefficient is \(5\) and y coefficient is \(-2\) So: \[Slope = - \frac{5}{-2} \implies \color{green}{Slope = \frac{5}{2}}\]

OpenStudy (anonymous):

I got it now, thanks everybody!

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

Generally after getting whole solution in one go everyone understands more easily.. Ha ha ha..